NAOJ GW Elog Logbook 3.2

We tried to calculate the beam profile but it seems that the beam divergence going in does not match the beam divergence going out.
i.e. approaching the beam waist, the beam shrinks ~ 100 um diameter per screw hole, but going past, it increases only ~ 50 um per screw hole.
We will look again for the source of this odd behaviour.

Logan, Michael
I showed Logan the use of the ATC clean room and the experiment setups we have there.
I showed and supervised the use of the Coherent Beam Master BM-3 UV beam profiler. Although I dropped and broke the included NG9 ND filter (0.5% transmission). The manual says to use below 5 mW and 200 W/mm^2 without the filter so it's fine in this case (3.8 mW).
We measured the beam profile from the s-polarized beam split off in the Taiwan/Korea experiment setups. I thought this beam was to be collimated at 2.3mm over some distance, but it doesn't maintain that size for long and turns out to have a waist of ~ 200 um radius about 30-40in from the PBS. So I also question a bit the Taiwan fiber SHG beam size, but we can look at that later. I tasked Logan with fitting the beam shape and redetermining the mode matching elements for the OPO.

File is being saved here after setting up new voltages
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Thu, May 2, 2024 12-12-22 PM.txt

The voltages for compensation can be found here:
foldername='C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\CompensationVoltage'
filename='voltagecompensating_0.02_0.002.txt'

The Jones matrix of the coating BB1EO3 was measured at 3 position 0,14,62 deg. This matrix alongwith BS was used to combine with the input polarization states to find the optimum voltages for compenstion. The desired voltage afterwards were saved in the format where there were range of other voltages obatined using the first set of voltage as locus to create a square of half edge 20mV with step size 2mV.
The coating was moved to 0 deg.
The voltage file was fed to the labview "Complete Integration-File Choice"
The measurement will be started after stable temperature of LC is attained. (warm up time is ususally 30mins)
The laser will be kept on till next week. So, kindly don't turn it off.
The voltages for compensation can be found here:
foldername='C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\CompensationVoltage'
filename='voltagecompensating_0.02_0.002.txt'
File is being saved here after setting up new voltages
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Thu, May 2, 2024 12-12-22 PM.txt
The measurement stopped because of windows update. I have paused windows update for next 5 weeks.
Started another one with new voltage set of radius 50mV, step 5mV. The condition on ell and azi was 5 degrees.The measurment is being saved with 1 average.
voltage file is here:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\CompensationVoltage\voltagecompensating_0.05_0.005.txt
find the measurment file here:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Sat, May 4, 2024 5-10-32 PM.txt
Couldn't transfer voltage file , a bit too huge
decreased condition to 20mV, 2mV and 2 degrees.
Also, I had made a mistake in conversion from rad to degree in the previous one , at one instance
voltage file:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\CompensationVoltage\voltagecompensating_0.02_0.002_2.txt
measurement file :
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Sat, May 4, 2024 6-05-39 PM.txt
Even though the voltages are being used from the file, it seems there could be a deviation, considering we did the calibration of LC 2 months ago. I can't see a zero crossing in azi and ell, also the file is a bit huge, so it will take longer than expected.
I am going to instead do scan using the voltage range estimated but without using the file, and doing first a rough scan, and then will later do a finer scan.
The voltage range estimated from characterization is 7-8 V for LC 1 and 0-25V for LC2 with condition of 2 degrees on azi and ell.
So, doing a scan from 6-9V with 0.2V step, and 0-25V, 0,2V step.
measurement file:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Sun, May 5, 2024 2-44-03 PM.txt
(could see zero crossing in labview, still need to plot, measurement finished by the time I finished writing this elog)
Did scans from
LC1, LC1step, LC2, LC2Step, averaging
0-25V, 0.2V, 0-25V, 0.2V
then observed and put a constraint on ell and azi of 2 deg
second scan
0-1V, 0.01V, 0-25V, 0.1V
15-25V, 0.1V, 0-25V, 0.1V
third scan
0-0.06V, 0.01V, 0-8V, 0.05V
15-25V, 0.01V, 0-5.2V, 0.01V
15-25V, 0.01V, 5.2-8V, 0.05V
0-0.06V, 0.01V, 0-4.3V, 0.01V
4th scan
0-0.04V, 0.005V, 0.83-1.2V, 0.001V
15-25V, 0.005V, 0.83-1.2V, 0.001V
5th scan
14.99-15.02, 0.001, 0-2V, 0.001V
24-25V, 0.001V, 0.95-0.994V, 0.001V

[Logan,Marc,Michael]
We took time to check all electronic connections and are now preparing a summary scheme.
During this activity, we also reconnected the 2 missing GRMC/CCFC cables.
We also replaced the broken CC PLL MON cable that was previously broken.
We could lock SHG in similar condition as before (large HOM) but could not lock GRMC despite relatively reasonable alignment.
THe issue is that DDS2 is not detecting the 500MHz clock signal while it is fine for DDS1 and 3.
We are now sending -20.5dBm clock to DDS2. To be further investigated

[Hugo,Logan,Marc]
We resetted the previous spyder environnement.
The unloadable libraries issues was solve by doing the appropriate pip install directly in the spyder console.
All usual python codes are working fine now.

There were issues to properly load python libraries on the pc so I resetted spyder settings to default.
Might be an issue for cv2?
[Hugo,Logan,Marc]
We resetted the previous spyder environnement.
The unloadable libraries issues was solve by doing the appropriate pip install directly in the spyder console.
All usual python codes are working fine now.

Logan, Michael
I introduced Logan to the squeezer clean room and mostly introduced FP cavity, FSR mode spectrum and PDH locking in the context of the green path.
I realigned the green mode cleaner/Mach Zender assembly to 92% mode matching. But the GRMC PD is not connected to the Cable Management Box. It's a bit hard to find in the giant mess of cables (seems to be unlabelled) so we will try next time.

The labview "Complete Integration for Camera" was modified to save data properly. The issue now was that the VI was too fast with averging of 1st order and there were same points being repeated in the dataset. Some changes were done as follows:
1. Added clear buffer for the averaging filter which is triggered when the number of iterations of averaging loop equals the averging order.
2. Added timer inside the averaging loop to make it a bit slower depending on what averging order you do. So, for different order you have different speed. Because the data saving was fine with averging order of 10, but was saving same data 4 times for averging order 1.

1 LCD à90° : Wed, Apr 24, 2024 3-40-14 PM.txt
1st 45° 2nd 0° : Wed, Apr 24, 2024 6-57-43 PM

Then, I made measurement with only 1 LCD of the polarization when we put a video on the LCD (video = all the 256 RGB value).
I did the same with 1 LCD oriented at 90°.
After doing this, I put the two different LCD and adapt my setup in order to get two different video, which allows you to combine the 256*256 different combinations.
All the setups seems to work.
Note : The camera is not synchronized with the 2 video make with a python.
1 LCD à90° : Wed, Apr 24, 2024 3-40-14 PM.txt
1st 45° 2nd 0° : Wed, Apr 24, 2024 6-57-43 PM

Mounted HWP on roator mount and measured the transmission using polcam. The measurement was taken at the following mentioned positions of the rotator mount.
The LC voltage was scanned from 0-3.5V with 0.1V step.
0deg:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Retarder\HWP\0 deg
Tue, Apr 23, 2024 4-45-43 PM.txt
38 deg:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Retarder\HWP\38 deg
Tue, Apr 23, 2024 5-10-09 PM.txt
The choice of position of rotator mount is random. The measurement is being taken to check if I can reconstruct rotation using the same method as being used for coating.

Mounted QWP on roator mount and measured the transmission using polcam. The measurement was taken at the following mentioned positions of the rotator mount.
The LC voltage was scanned from 0-3.5V with 0.1V step.
0deg:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Retarder\QWP\0 deg
Tue, Apr 23, 2024 11-43-44 AM.txt
62 deg:
C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Retarder\QWP\62 deg
Tue, Apr 23, 2024 11-57-20 AM.txt

The pressure at 8:00 was 4.7*10-4 Pa. Looking at previous logs, the final pressure is consistent with past values.
I plan to start cooling on Thursday.
I also checked the valves for water supply to compressor are open. So I might not have to worry about cleaning up the water lines. I heard the water is supplied by ATC, but do I need to turn on the water chiller?

It seems there could be some issue with previous data, as there is some issue with reconstructing rotation of the coating. I took measurement at several position of coating today. The coating BB1EO3 is mounted on a rotator and hence the rotator is rotated at 0 deg, 14deg and 62deg. The voltage scanned was from 0 to 3.5V with 0.1V resolution and averaging of 10th order is implemented.
Also the rotator mount was changed. previoulsy we were using rotator alone and hence it was quite difficult to have finely tuned alignment between the incident and reflected beam. The rotator is now mounted like this and now we could perfectly align the reflected beam with incident beam after the BS
path = "C:/Users/atama/OneDrive/LC-Experiment/Measurement Data/"
0 deg
14 deg

I turned on the TMP at 12:54. The pressure at that time was 3.58*101 Pa.
The pressure at 13:05 was 8.1*10-2 Pa.
I will continue pumping for a while (until Monday).
Note: The TMP is in "local" mode and can only be operated using the front control panel.

The voltage of 0-3.5V was chosen in contrast to previous choice of 0-5V or 0-25V. It is because most of changes induced in azimuth and ellipticity due to LC is around this range of voltage.

An old elog that I forgot to post and was in my drafts.
The LC were scanned from 0-3..5V with 0.1V resolution
Measurements were taken at coating position of 0 and 28deg
All the measurement were saved with the the labview where we can even save std on each parameter.
Input_foldername=os.path.join(path, 'Polarization states/20240401')
input_filename= 'Mon, Apr 1, 2024 9-59-08 AM.txt'
for 0 deg
Output_foldername = os.path.join(path, 'Coating measurement/BB1-E03/20240401/0 deg')
output_filename ='Mon, Apr 1, 2024 3-52-42 PM.txt'
for 28 deg
Output_foldername = os.path.join(path, 'Coating measurement/BB1-E03/20240401/28 deg')
output_filename ='Mon, Apr 1, 2024 9-25-18 AM.txt'
The voltage of 0-3.5V was chosen in contrast to previous choice of 0-5V or 0-25V. It is because most of changes induced in azimuth and ellipticity due to LC is around this range of voltage.

4 different video for : fast or low video and different power of the initial beam.
We tried to calculate the beam profile but it seems that the beam divergence going in does not match the beam divergence going out.
i.e. approaching the beam waist, the beam shrinks ~ 100 um diameter per screw hole, but going past, it increases only ~ 50 um per screw hole.
We will look again for the source of this odd behaviour.